OpenStudy (anonymous):

Solve the equation 2(3x + 7) = -22.

OpenStudy (anonymous):

first divide both sides by 2 than take minus 7 on both times divide that by 3 and you got your answer

OpenStudy (anonymous):

Just do the opposite of PEMDAS. 2(3x + 7) = -22 3x + 7 = -11 3x = -18 x = -6

OpenStudy (anonymous):

Thank youuuu !

OpenStudy (anonymous):

don't do the math for them... than they never learn...

OpenStudy (anonymous):

2(3x+7)=-22 6x+14=-22 -14 -14 ___________________ 6x= -36 x=-6 First distribute everything within the parenthesis, then work the rest algebraically